OpenStudy (anonymous):

evaluate the integral of sec^2(x)/sqrt(1+tan(x) Show work?

OpenStudy (xapproachesinfinity):

@Alexkwkwl still need help

OpenStudy (anonymous):

Yes please

OpenStudy (xapproachesinfinity):

So what are your thoughts

OpenStudy (xapproachesinfinity):

let's just make sure we have the right integral \[\int \frac{\sec^2x}{\sqrt{1+\tan x}}dx\]

OpenStudy (xapproachesinfinity):

is this what you have

OpenStudy (anonymous):

Yes

OpenStudy (xapproachesinfinity):

how about a sub in there?

OpenStudy (anonymous):

A sub?

OpenStudy (xapproachesinfinity):

substitution

OpenStudy (anonymous):

No

OpenStudy (xapproachesinfinity):

why not?

OpenStudy (anonymous):

Oh you mean like 1/cos

OpenStudy (xapproachesinfinity):

this sub works good \[u=1+\tan x\]

OpenStudy (xapproachesinfinity):

\[du=\sec^2x dx \]

OpenStudy (xapproachesinfinity):

you can see it now?

OpenStudy (anonymous):

Wait And then what do you do

OpenStudy (xapproachesinfinity):

you substitute \(\sec^2x dx\) for \(du\)

OpenStudy (anonymous):

Yes i know that

OpenStudy (xapproachesinfinity):

and \(1+\tan x \) for u

OpenStudy (xapproachesinfinity):

then work the integral with u

OpenStudy (xapproachesinfinity):

so?

OpenStudy (xapproachesinfinity):

this is what you get \[\int \frac{du}{ \sqrt{u}}\]

OpenStudy (xapproachesinfinity):

you should be able to integrate this one easily

OpenStudy (anonymous):

2?

OpenStudy (xapproachesinfinity):

what you mean by two? remember indefinite integral gives a family of functions meaning an anti-derivative with a constant C

OpenStudy (xapproachesinfinity):

you should not get a numerical value

OpenStudy (anonymous):

2 times sqrtu +c

OpenStudy (xapproachesinfinity):

yes! now revert back to 1+tanx

OpenStudy (xapproachesinfinity):

remember that we set u=1+tanx

OpenStudy (anonymous):

2(1+Tanx) +C

OpenStudy (xapproachesinfinity):

no

OpenStudy (xapproachesinfinity):

you forgot something

OpenStudy (anonymous):

Oh the sqrt lol duh

OpenStudy (xapproachesinfinity):

yeah

OpenStudy (anonymous):

That's it??

OpenStudy (xapproachesinfinity):

yep

OpenStudy (anonymous):

Omg yay thank you so much

OpenStudy (xapproachesinfinity):

yeah you doing great :) you are welcome